OpenStudy (konradzuse):

Use the integral test to determine the convergence or divergence of the series.

OpenStudy (konradzuse):

\[\sum_{}^{\infty} \frac{\ln(n)}{n^{10}}\]

OpenStudy (konradzuse):

so if it's an integral iut's going to be .

OpenStudy (konradzuse):

\[\int\limits_{2}^{\infty} \ln(x)/x^{10}\]

OpenStudy (konradzuse):

\[\int\limits_{2}^{\infty} \ln(x) * \frac{1}{x^{10}}\]

OpenStudy (anonymous):

solve for it

OpenStudy (konradzuse):

u = ln(x) du = 1/x dx

OpenStudy (konradzuse):

dv = 1/x^10 v = -1/9x^9

OpenStudy (konradzuse):

ok now integration by parts.

OpenStudy (konradzuse):

\[-\frac{\ln(x)}{9x^9} +\frac{1}{9} \int\limits \frac{1}{x^{10}}dx\]

OpenStudy (konradzuse):

\[-\frac{\ln(x)}{9x^9} +\frac{1}{9} * -\frac{1}{9x^{9}} + c\]

OpenStudy (anonymous):

evaluate it at the limits

OpenStudy (anonymous):